# Data-Driven High-Dimensional Statistical Inference with Generative Models

### v1

@@ Abstract
Crucial to many measurements at the LHC is the use of correlated multi-dimensional information to distinguish rare processes from large backgrounds, which is complicated by the poor modeling of many of the crucial backgrounds in Monte Carlo simulations. In this work, we introduce HI-SIGMA, a method to perform unbinned high-dimensional statistical inference with data-driven background distributions. In contradistinction to many applications of Simulation Based Inference in High Energy Physics, HI-SIGMA relies on generative ML models, rather than classifiers, to learn the signal and background distributions in the high-dimensional space. These ML models allow for efficient, interpretable inference while also incorporating model errors and other sources of systematic uncertainties. We showcase this methodology on a simplified version of a di-Higgs measurement in the $bb\gamma\gamma$ final state, where the di-photon resonance allows for efficient background interpolation from sidebands into the signal region. We demonstrate that HI-SIGMA provides improved sensitivity as compared to standard classifier-based methods, and that systematic uncertainties can be straightforwardly incorporated by extending methods which have been used for histogram based analyses.
@@ 1 Introduction
The most common usage of classifier in analysis is by placing a cut on the classifier score to reject background. Events passing these selections are then fit in some discriminating feature, such as the mass of a resonance candidate, to extract the relevant parameter of interest. A more optimal, and increasingly common, alternate strategy is to use a binned classifier score as the final fit observable rather than a cut variable. This approach results in improved sensitivity because it makes better use of the information learned by the classifier; cuts throw away some signal events from statistical analysis, and discard information about the relative signal-versus-background likelihood of the remaining events. However, it can be difficult to estimate the background in a classifier score fit analysis because multivariate correlations among all classification features must be accounted for. For many analyses, known mismodelings of backgrounds in simulation necessitate data-driven background estimates, which can be difficult in the multivariate case. Binned classifier score fits can also suffer from a lack of interpretability; the final fit observable is no longer physical and represents a collapse of a complex multivariate space to a single observable. This obscures the intuition of analyzers and make it difficult to spot potential mismodelings. Such questions of interpretability may come to the forefront should any such analysis report statistical evidence of beyond standard model physics.
@@
Additionally, the use of binned classifiers are close to optimal only for analyses of a single parameter of interest with linear parameter dependenceBrehmer:2018eca , and therefore will require undesirable tradeoffs in the case of signal inference or multi-parameter extraction, such as measurements constraining multiple effective field theory operators. A class of analysis techniques, called simulation based inference (SBI) has been proposed to optimally perform inference in this case Brehmer:2018kdj ; Brehmer:2018eca ; Cranmer_sbi_review ; Ghosh_jrjc_proc . Nearly all SBI proposals in particle physics are based on classifiers, making use of the property that appropriately trained neural network classifiers learn the likelihood ratio between their target classes Cranmer:2015bka . These SBI methods therefore train classifiers on simulation to learn the likelihood ratio between different hypotheses under consideration, and directly use the learned likelihood ratios to perform statistical inference. The first application of this technique to a real LHC analysis has recently been performed by the ATLAS Collaboration ATLAS_SBI_methods ; ATLAS_SBI_measurement . SBI methods allow for optimal multivariate inference, but have a heavy reliance on simulation, which hinders their applicability to final states which are known to require data-driven modeling.
@@
In this work we introduce a new approach, High Dimensional Statistical Inference with Generative Model AI (HI-SIGMA), which can perform optimal multivariate statistical inference with data-driven background modeling. We propose to use machine learning models to learn the multi-dimensional probability densities of signal and background, and perform statistical inference directly on these learned probability densities, rather than relying on classifiers. We consider a resonance analysis, in which the signal is known to be localized in some signal region (SR), with sidebands on either side dominated by background events. Borrowing from techniques previously employed in anomaly detection Hallin:2021wme , we use a generative model to learn the conditional multi-dimensional probability density of the background from the sideband regions, and interpolate it into the signal region. The background probability density in the single resonant variable is estimated with a standard parametric functional form. These two pieces and then combined for a full estimate of the multi-dimensional background probability density. The multi-dimensional signal probability density is learned with Monte Carlo simulations. These two densities are then used to construct the likelihood function of the signal plus background hypothesis, in an analogous fashion to standard histogram-based template analyses. A crucial challenge in employing ML-based density estimators within statistical inference is the appropriate assessment of uncertainties to the learned density estimates. We show how techniques which are standard practice in uncertainty quantification in histogram-template-based density estimates can be extended to quantify the uncertainty on these ML-based density estimators.
@@
This paper is organized as follows. In Section 2, we outline the HI-SIGMA method and in Section 3 we discuss the incorporation of systematic uncertainties. We then demonstrate an application of HI-SIGMA to a di-Higgs measurement in Section 4. We discuss the various pros and cons of the HI-SIGMA method as compared to commonly employed analysis strategies in Section 5, and conclude in Section 6.
@@ 2 The HI-SIGMA Method
Recently, Ref. Sluijter:2025isc compared generative and classifier-based approaches to SBI in a particle physics context and performed coverage studies. On both a toy Gaussian example and a mock $H\to\tau\tau$ measurement, it found both approaches to have proper coverage, but the generative approach led to slightly larger uncertainties. However, Ref. Sluijter:2025isc trained the generative models using only simulation, and did not consider the data-driven estimation of the background density featured in the HI-SIGMA approach. Ref. Sluijter:2025isc also performed its studies on only a two-dimensional feature set, rather than the 4+1 dimensional example we showcase in our $bb\gamma\gamma$ example.
@@ 2.1 Formalism
where $P_{k}(m)$ is a one-dimensional the probability density of the resonance variable. This one-dimensional background $m$ density can be straightforwardly estimated from the data using the well-tested parametric functional forms used in countless collider analyses. [^1] Alternatively, Gaussian process regression could be used if a suitable parametric form cannot be foundFrate:2017mai ; Gandrakota:2022wyl .
@@ 3 Incorporating systematics
Each piece of our likelihood function will have systematic uncertainties which need to be incorporated for reliable inference. Uncertainties on the parametric shape for the signal and background mass distributions can be handled in typical fashion. The background shape nuisance parameters are left freely floating in the fit, since they are constrained by the sideband data, and profiled over; while signal shape parameters and uncertainties are usually determined from simulation. Uncertainties affecting the multidimensional feature density distributions are more important and require novel strategies. We consider two different types of uncertainty affecting the learned densities.
@@ 3.1 Uncertainty from Finite Training Data
The first uncertainties we consider are statistical uncertainties arising from finite events, either in data sideband regions or Monte Carlo simulations, available for training the generative models. To account for these uncertainties we bootstrap our set of training events into sub-samples, and train a separate generative model for each sub-sample. Statistical inference can then be performed with these separate bootstrapped models, and their likelihood curves combined in some fashion to determine the overall best fit and uncertainty. One option to combine the different likelihood curves would be to use discrete profiling Dauncey:2014xga , which has been used in high energy physics to account for the choice of parametric functional form. In discrete profiling, for each value of the parameter of interest (p.o.i.) the overall log-likelihood is defined as the minimum of the log-likelihoods of the considered models. This allows the data to ’constrain’ the discrete nuisance defining the choice of model; if the data prefers one model such that its log-likelihood is less than all others within the ±$\pm 1\sigma$ confidence interval of the p.o.i., then discrete profiling will not increase the size of the confidence interval. In the case of alternate functional forms with low numbers of parameters, such a preference may be realistic, however in the case of profiling over different generative models a more conservative approach may be warranted.
In the HI-SIGMA application of this method, the models being profiled over each have the same ’functional form’ but have a very large number of free parameters. Neural networks with large numbers of parameters are also known to essentially never reach their global minima, which is nearly always an overfitted model. Different models may therefore end up in quite distinct minima. However, when fitting large datasets, of say $O(10^{4})$ events, a $O(10^{-3})$ difference in the log-likelihood of the different models will result in a $O(10)$ shift in the best fit minima, meaning discrete profiling will depend only on the ’best’ model within the ±$\pm 3\sigma$ confidence interval. We therefore propose a more conservative approach, in which we define the final result as an envelope of the $\Delta NLL$ curves of the different bootstrapped models.
Thus, if we have $D_{b}$ bootstrapped SB samples for background estimation and $D_{s}$ bootstrapped MC samples for signal estimation, we have $d=1,\dots,D_{b}\times D_{s}$ profiled likelihoods for each $s$, $\mathcal{L}_{d}(s,\hat{\hat{\nu}})$, and we obtain the confidence intervals by combining the “minimum substracted models” into the modified test statistic
@@
Note, in our implementation in Section 4, for improved stability we take each bootstrapped model to be an ensemble of five generative models trained on the same dataset with different initial seeds. The log-probabilities of each individual model are averaged to produce the final density estimate.
@@ 3.2 Shape uncertainties
An additional important uncertainty to consider are variations of the feature distributions from known sources of systematic uncertainty. In histogram-based analyses, these are accounted for by producing an ‘up’ and ‘down’ variation of the relevant distribution, corresponding to ±$\pm 1\sigma$ variations of the systematic uncertainty. A nuisance parameter, $\nu$, is included into the fit which captures these systematic variations. We define $\nu=0$ as the nominal distribution, whereas $\nu=\pm 1$ represents the up/down variations. A Gaussian constraint is added to the likelihood to penalize variations away from $\nu=0$. For intermediate values of $\nu$, an interpolation scheme is used to produce an estimate of the density in between the nominal and up/down variations (see Ref. Cranmer:2014lly for a review).
@@
$w_{k}(\vec{x}^{\prime},m,\nu)=\begin{cases}\left(\frac{P_{k}(\vec{x}^{\prime}|m,\nu=1)}{P_{k}(\vec{x}^{\prime}|m,\nu=0)}\right)^{\nu}&\text{ if $\nu>1$}\\ 1+\sum_{i=1}^{6}c_{ki}(\vec{x}^{\prime},m)\nu^{i}&\text{ if $|\nu|\leq 1$}\\ \left(\frac{P_{k}(\vec{x}^{\prime}|m,\nu=1)}{P_{k}(\vec{x}^{\prime}|m,\nu=0)}\right)^{-\nu}&\text{ if $\nu<-1$}\\ \end{cases}$ (8)
and $P_{k}(\vec{x}^{\prime}|m,\nu=0)$ ($P_{k}(\vec{x}^{\prime}|m,\nu=\pm 1)$) is the nominal (±$\pm 1\sigma$) model as obtained from averaging over the ensemble of models obtained with the nominal (±$\pm 1\sigma$) systematic variation of the MC samples, and $c_{ki}(\vec{x}^{\prime},m)$ are chosen so that the function and its derivatives are continuous at $\nu=\pm 1$. Note that the interpolation coefficients, $c_{ki}(\vec{x}^{\prime},m)$, are determined uniquely for each data point, because all points must have continuous density estimates. This an extension of the binned case in which the coefficients are determined separately for each bin.
The weights are incorporated into the likelihood model and $\nu$ is profiled over. This approximation allows to include in principle several systematic sources which are assumed to be independent while keeping the number of trained models relatively under control.

### v2

@@ Abstract
Crucial to many measurements at the LHC is the use of correlated multi-dimensional information to distinguish rare processes from large backgrounds, which is complicated by the poor modeling of many of the crucial backgrounds in Monte Carlo simulations. In this work, we introduce HI-SIGMA, a method to perform unbinned high-dimensional statistical inference with data-driven background distributions. In contradistinction to many applications of Simulation Based Inference in High Energy Physics, HI-SIGMA relies on generative ML models, rather than classifiers, to learn the signal and background distributions in the high-dimensional space. These ML models allow for interpretable inference while also incorporating model errors and other sources of systematic uncertainties. We showcase this methodology on a simplified version of a di-Higgs measurement in the $bb\gamma\gamma$ final state, where the di-photon resonance allows for background interpolation from sidebands into the signal region. We demonstrate that HI-SIGMA provides improved sensitivity as compared to standard classifier-based methods, and that systematic uncertainties can be straightforwardly incorporated by extending methods which have been used for histogram based analyses.
@@ 1 Introduction
The most common usage of classifier in analysis is by placing a cut on the classifier score to reject background. Events passing these selections are then fit in some discriminating feature, such as the mass of a resonance candidate, to extract the relevant parameter of interest. A more optimal, and increasingly common, alternate strategy is to use a binned classifier score as the final fit observable rather than a cut variable. This approach results in improved sensitivity because it makes better use of the information learned by the classifier; cuts throw away some signal events from statistical analysis, and discard information about the relative signal-versus-background likelihood of the remaining events. The usage of binned classifier scores can be understood as a natural extension of the Matrix Element Method Kondo:1988yd; Kondo:1991dw, or, for the case of linear parameter dependence, the ‘optimal observables’ framework Diehl:1993br. The difference being that the Matrix Element Method uses first-principles parton-level likelihood ratios in conjunction with smearing methods to approximate non-perturbative and detector effects, a trained binary classifier learns the likelihood ratio directly from the reconstructed observables.
Despite the power of these approaches, they also have challenges. Most importantly, it can be difficult to estimate the background in a classifier score fit analysis because multivariate correlations among all classification features must be accounted for. For many analyses, known mismodelings of backgrounds in simulation necessitate data-driven background estimates, which can be difficult in the multivariate case. Binned classifier score fits can also suffer from a lack of interpretability; the final fit observable is no longer physical and represents a collapse of a complex multivariate space to a single observable. This obscures the intuition of analyzers and make it difficult to spot potential mismodelings. Such questions of interpretability may come to the forefront should any such analysis report statistical evidence of beyond standard model physics.
@@
Additionally, the use of binned classifiers are close to optimal only for analyses of a single parameter of interest with linear parameter dependenceBrehmer:2018eca, and therefore will require undesirable tradeoffs in the case of signal inference or multi-parameter extraction, such as measurements constraining multiple effective field theory operators. A class of analysis techniques, called simulation based inference (SBI) has been proposed to optimally perform inference in this case Brehmer:2018kdj; Brehmer:2018eca; Cranmer_sbi_review.
Nearly all SBI proposals in particle physics are based on classifiers, making use of the property that appropriately trained neural network classifiers learn the likelihood ratio between their target classes Cranmer:2015bka. These SBI methods therefore train classifiers on simulation to learn the likelihood ratio between different hypotheses under consideration, and directly use the learned likelihood ratios to perform statistical inference Ghosh_jrjc_proc; GomezAmbrosio:2022mpm; Bahl:2021dnc; Barrue:2023ysk; Schofbeck:2024zjo; Chai:2024zyl; nsbi_dihiggs; Benato:2025rgo. The first application of this technique to a real LHC analysis has recently been performed by the ATLAS Collaboration ATLAS_SBI_methods; ATLAS_SBI_measurement. SBI methods allow for optimal multivariate inference, but have a heavy reliance on simulation, which hinders their applicability to final states which are known to require data-driven modeling.
@@
In this work we introduce a new approach, High Dimensional Statistical Inference with Generative Model AI (HI-SIGMA), which can perform optimal multivariate statistical inference with data-driven background modeling. We propose to use machine learning models to learn the multi-dimensional probability densities of signal and background, and perform statistical inference directly on these learned probability densities, rather than relying on classifiers. We consider a resonance analysis, in which the signal is known to be localized in some signal region (SR), with sidebands on either side dominated by background events. Borrowing from techniques previously employed in anomaly detection Hallin:2021wme, we use a generative model to learn the conditional multi-dimensional probability density of the background from the sideband regions, and interpolate it into the signal region. The background probability density in the single resonant variable is estimated with a standard parametric functional form. These two pieces and then combined for a full estimate of the multi-dimensional background probability density. The multi-dimensional signal probability density is learned from Monte Carlo simulations. These two densities are then used to construct the likelihood function of the signal plus background hypothesis, in an analogous fashion to standard histogram-based template analyses. A crucial challenge in employing ML-based density estimators within statistical inference is the appropriate assessment of uncertainties to the learned density estimates. We show how techniques which are standard practice in uncertainty quantification in histogram-template-based density estimates can be extended to quantify the uncertainty on these ML-based density estimators.
@@
This paper is organized as follows. In Section 2, we outline the HI-SIGMA method and in Section 3 we discuss the incorporation of systematic uncertainties. We then detail an application of HI-SIGMA to a di-Higgs measurement in Section 4 and show the results in Section 5. We discuss the various pros and cons of the HI-SIGMA method as compared to commonly employed analysis strategies in Section 6, and conclude in Section 7.
@@ 2 The HI-SIGMA Method
Recently, Ref. Sluijter:2025isc compared generative and classifier-based approaches to SBI in a particle physics context and performed coverage studies. On both a toy Gaussian example and a mock $H\to\tau\tau$ measurement, it found both approaches to have proper coverage, but the generative approach led to slightly larger uncertainties. However, Ref. Sluijter:2025isc trained the generative models using only simulation, and did not consider the data-driven estimation of the background density featured in the HI-SIGMA approach. Also, due to the numerical requirements of their large suite of pseudoexperiments, Ref. Sluijter:2025isc performed its studies on only a two-dimensional feature set, rather than the 4+1 dimensional example we showcase in our $bb\gamma\gamma$ example. As we detail below, we find that the performance of HI-SIGMA is dependent on the dimensionality of feature space. It can more difficult to accurately model the probability density of higher dimensional feature spaces, but they offer potentially more information to perform inference with.
@@ 2.1 Formalism
where $P_{k}(m)$ is a one-dimensional the probability density of the resonance variable. This one-dimensional background $m$ density can be straightforwardly estimated from the data using the well-tested parametric functional forms used in countless collider analyses. [^1] Alternatively, Gaussian process regression could be used if a suitable parametric form cannot be foundFrate:2017mai; Gandrakota:2022wyl; Barr:2025lba.
@@
Although we consider only a simple one parameter measurement for this first study, the formalism can be easily extended to inference with multiple parameters of interest $\vec{\theta}$. Often in particle physics the likelihood has the form of a mixture model where the parameters of interest control the relative contributions of the different processes. This means the conditional likelihood of the different processes can be learned separately and independently of $\vec{\theta}$.
If the dependence of the likelihood in terms of the different parameters is not known a priori, MC simulations need to be generated for different values of $\vec{\theta}$ and the generative model can be trained to learn the likelihood conditional on $\vec{\theta}$, $P(\vec{x}^{\prime}|m,\vec{\theta})$. Since we are learning the conditional likelihood, as in most SBI methods, the particular choice of prior distribution over parameters of interest is not important as long as the relevant region in parameter space is sufficiently covered. Having estimated $P_{s}(\vec{x}^{\prime}|m,\vec{\theta})$ and $P_{s}(m|\nu_{s},\vec{\theta})$, we can perform multi-dimensional fits on $\vec{\theta}$ using any of the standard fitting tools. Thus, even if the problem becomes numerically more challenging, it is not conceptually different.
@@ 3 Incorporating systematics
Each piece of our likelihood function will have systematic uncertainties which need to be incorporated for reliable inference. Uncertainties on the parametric shape for the signal and background mass distributions can be handled in typical fashion. The background shape nuisance parameters are left freely floating in the fit, since they are constrained by the sideband data, and profiled over; while signal shape parameters and uncertainties are usually determined from simulation. Uncertainties affecting the multi-dimensional feature density distributions are more important and require novel strategies.
The increased information available when fitting in high dimensions can break degeneracies between shape uncertainties and the parameter of interest, reducing the impact of systematic uncertainties. However, this increased precision requires a corresponding increase in the rigor of the systematic uncertainty parameterization to ensure proper analysis coverage.
We consider two different types of uncertainty affecting the learned densities.
@@ 3.1 Uncertainty from Finite Training Data
The first uncertainties we consider are statistical uncertainties arising from finite events, either in data sideband regions or Monte Carlo simulations, available for training the generative models. To account for these uncertainties we bootstrap our set of training events into sub-samples, and train a separate generative model for each sub-sample. Statistical inference can then be performed with these separate bootstrapped models, and their likelihood curves combined in some fashion to determine the overall best fit and uncertainty. One option to combine the different likelihood curves would be to use discrete profiling Dauncey:2014xga, which has been used in high energy physics to account for the choice of parametric functional form. In discrete profiling, for each value of the parameter of interest (p.o.i.) the overall log-likelihood is defined as the minimum of the log-likelihoods of the considered models. This allows the data to ‘constrain’ the discrete nuisance defining the choice of model; if the data prefers one model such that its log-likelihood is less than all others within the $\pm 1\sigma$ confidence interval of the p.o.i., then discrete profiling will not increase the size of the confidence interval. In the case of alternate functional forms with low numbers of parameters, such a preference may be realistic, however in the case of profiling over different generative models a more conservative approach may be warranted.
In the HI-SIGMA application of this method, the models being profiled over each have the same ‘functional form’ but have a very large number of free parameters. Neural networks with large numbers of parameters are also known to essentially never reach their global minima, which is nearly always an overfitted model. Different models may therefore end up in quite distinct minima. However, when fitting large datasets, of say $O(10^{4})$ events, a $O(10^{-3})$ difference in the log-likelihood of the different models will result in a $O(10)$ shift in the best fit minima, meaning discrete profiling will depend only on the ‘best’ model within the $\pm 3\sigma$ confidence interval. We therefore propose a more conservative approach, in which we define the final result as an envelope of the $\Delta NLL$ curves of the different bootstrapped models.
To reduce the impact of outlier models which fit the data poorly, we require that the log-likelihood difference between the fit performed with a given model and one performed the ‘best’ model (the fit with the maximum likelihood) be less than 50, to be included in the envelope. The threshold of 50 was not optimized. It represents a ${\sim}7\sigma$ preference of the data for one model versus the other. Empirically, this requirement allows the rejection of extreme outliers while preserving the conservative approach of the uncertainty.
[…]
@@
Note, in our implementation in Section 4, for improved stability we take each bootstrapped model to be an ensemble of ten generative models trained on the same dataset with different initial seeds. To further prevent overfitting, we divide the boostrapped sample in training and validation samples, and stop the training if the validation loss has not improved after a certain number of epochs. The final model is taken as the model from the epoch with the lowest validation loss. The log-probabilities of each individual model are averaged to produce the final density estimate. We find that this procedure of early stopping and ensembling reduces model variability between the different bootstraps.
@@ 3.2 Shape uncertainties
An additional important uncertainty to consider are variations of the feature distributions from known sources of systematic uncertainty. In histogram-based analyses, these are accounted for by producing an ‘up’ and ‘down’ variation of the relevant distribution, corresponding to $\pm 1\sigma$ variations of the systematic uncertainty. A nuisance parameter, $\nu$, is included into the fit which captures these systematic variations. We define $\nu=0$ as the nominal distribution, whereas $\nu=\pm 1$ represents the up/down variations. A Gaussian constraint is added to the likelihood to penalize variations away from $\nu=0$. For intermediate values of $\nu$, an interpolation scheme is used to produce an estimate of the density in between the nominal and up/down variations (see Ref. Cranmer:2014lly for a review).
@@
[…]
and $P_{k}(\vec{x}^{\prime}|m,\nu=0)$ ($P_{k}(\vec{x}^{\prime}|m,\nu=\pm 1)$) is the nominal ($\pm 1\sigma$) model as obtained from averaging over the ensemble of models obtained with the nominal ($\pm 1\sigma$) systematic variation of the MC samples, and $c_{ki}(\vec{x}^{\prime},m)$ are chosen so that the function and its first two derivatives are continuous at $\nu=\pm 1$. The continuity condition defines a set of six linear equations which can be used to solve explicitly for the interpolation coefficients $c_{ki}(\vec{x}^{\prime},m)$:
$\begin{split}1+\sum_{i=1}^{6}c_{ki}(\vec{x}^{\prime},m)(\pm 1)^{i}&=\frac{P_{k}(\vec{x}^{\prime}|m,\nu=\pm 1)}{P_{k}(\vec{x}^{\prime}|m,\nu=0)}\,,\\ \sum_{i=1}^{6}c_{ki}(\vec{x}^{\prime},m)i(\pm 1)^{i-1}&=\pm\frac{P_{k}(\vec{x}^{\prime}|m,\nu=\pm 1)}{P_{k}(\vec{x}^{\prime}|m,\nu=0)}\ln\frac{P_{k}(\vec{x}^{\prime}|m,\nu=\pm 1)}{P_{k}(\vec{x}^{\prime}|m,\nu=0)}\,,\\ \sum_{i=2}^{6}c_{ki}(\vec{x}^{\prime},m)i(i-1)(\pm 1)^{i-2}&=\frac{P_{k}(\vec{x}^{\prime}|m,\nu=\pm 1)}{P_{k}(\vec{x}^{\prime}|m,\nu=0)}\left(\ln\frac{P_{k}(\vec{x}^{\prime}|m,\nu=\pm 1)}{P_{k}(\vec{x}^{\prime}|m,\nu=0)}\right)^{2}\,.\end{split}$
Note that these interpolation coefficients, $c_{ki}(\vec{x}^{\prime},m)$, are determined uniquely for each data point, because all points must have continuous density estimates. This an extension of the binned case in which the coefficients are determined separately for each bin. The weights are incorporated into the likelihood model and $\nu$ is profiled over. This approximation allows to include in principle several systematic sources which are assumed to be independent while keeping the number of trained models relatively under control.
Although the flexibility of Eq.9 permits the modeling of complex shape uncertainties, the parameterization can be too flexible in certain cases. In particular, we empirically found that the coefficients overfit what turns out to be a simpler dependence on $\nu$ for the example considered in Section 4. We thus consider a less complex parameterization of the weights,
$w_{k}(\vec{x}^{\prime},m,\nu)=1+\sum_{i=1}^{2}c_{ki}(\vec{x}^{\prime},m)\nu^{i}$ (9)
with no piecewise splitting and only two learnable coefficients, which can be fitted by evaluating the weights at $\nu=\pm 1$. We empirically validate this choice through a closure test, detailed in Section 4. Although in this work we do not consider the full form of Eq. 8, it is useful to understand 9 as its simplification and we expect the former to be necessary for more complex shape uncertainties.